Cu Determination (5 Points: 5 Points titration, 10 This is for an analytical chemistry. Please show all work and excel sheets.
Answer #1

EDTA is a chelating agent and chelates metal ions, i.e. attaches metal ions to itself.

For the Cu titration,

taking the average value of the 7 trials for the volume of EDTA consumed, we get:

V EDTA (avg) = 16.90 mL

Assume that the tap water has a Cu2+ conc. of x moles/L

Volume of aliquot = 25 mL = 0.025 L

Moles of Cu2+ in this aliquot = 0.025*x moles

These many moles of EDTA will also be required for titration.

Lets say the standard EDTA solution used in this experiment has EDTA conc. of 'y' moles/L

Volume of EDTA consumed = 16.90 mL = 0.0169 L

Thus, moles of EDTA used = 0.0169*y moles

Equating we get:
0.0169*y = 0.025*x

---------------------------------

For the Pb2+ titration:
Lets assume that tap water has Pb2+ conc of 'x' moles/L, standard EDTA solution has conc 'y' moles/L and standard Mg2+ solution has conc. 'z' moles/L.

Moles of EDTA supplied = Volume*conc = 0.025*y

Moles of Pb2+ present = 0.050*x

Moles of EDTA left after the reaction of EDTA with Pb2+ = 0.025*y - 0.050*x

These many moles of Mg2+ are supplied from outside to react with EDTA.

Calculating the average volume of Mg solution used in titration from the 6 trials, we get;

V Mg (average) = 19.26 mL

Thus, moles of Mg2+ = 0.01926*z

Thus, equating we get:

0.01926 = 0.025*y - 0.050*x

Putting 'y' and 'z', we can calculate 'x'